OpenStudy (anonymous):

the 9th roots of unity ? the 9th roots of unity ? @Mathematics

OpenStudy (anonymous):

What about them?

OpenStudy (anonymous):

x ^ 9 = 1 right ?

OpenStudy (anonymous):

(x^3)^3-1^3 =0 (x^3-1)(x^6+x^3+1)=0 x=-1,-1,-1 ??? how do we find them

OpenStudy (anonymous):

Sorry I'm not of much assistance, but perhaps this helps? http://paul-mccarthy.us/Misc/ProvingRelations.htm